# Controlling Impact Noise Propagation With Ceramic Metatiles

### v1

@@ Abstract
We examine the application of metamaterials for impact noise insulation in tiled floors. A ceramic metatile is developed which makes use of phononic metamaterials and optimized joint configurations. First, we optimize the geometrical and material parameters of the proposed metatile, which is composed of small ceramic tiles connected by silicon joints, in order to reduce longitudinal and flexural wave propagation on tiled floors. A bandgap is achieved that effectively mitigates the transmission of impact noise through periodic structural pathways. In both experimental and numerical tests, it is demonstrated that the integration of silicon joints inside the ceramic metatile improves the acoustic insulation performance, as measured by the reduction of impact noise levels across a wide range of low frequencies. The findings highlight the potential of metamaterials in architectural acoustics, offering innovative solutions for sound control in tiled environments.
I Introduction
@@
One of the most significant factors in building design is to provide a comfortable environment for its residents. Several elements influence a home’s degree of comfort, including temperature, humidity, noise, air quality and hygiene. Neglecting such living conditions will have a negative impact on household comfort, e.g., noise generated by footsteps and noise pollution from surrounding buildings, temperature rise due to heat absorbed by buildings during the daytime, inadequate sanitation, and poor indoor air quality [1, 2]. Among these living conditions, noise control is a critical aspect of architectural design and urban planning, with far-reaching implications for physical and mental health. Reducing noise in enclosed spaces is a long-standing objective in architectural acoustics, especially in contexts like tiled rooms, where sound reflections and transmissions are enhanced by hard, reflective surfaces [3]. In general, there are numerous ways of controlling noise annoyance both within and outside buildings. Typical noise mitigation methods, such as the use of porous absorbers or heavy insulation, are often encountered with aesthetic, functional, or structural constraints. Recent advances in metamaterials [4, 5, 6, 7, 8, 9], including phononic crystals and acoustic metasurfaces, have opened new avenues for sound control in buildings, enabling the manipulation of wave propagation using compact and tunable designs [3, 10]. Thus, the solution to improve the comfort of noise insulation in building floors should be based on phononic metamaterials in one part and institutive materials in the other part.
Phononic metamaterials, which consist of periodic arrangements of scatterers and resonators with contrasting acoustic properties, are particularly promising due to their ability to create low frequency band gaps that prevent propagation of specific wave modes [11, 12, 13]. These bandgaps arise from Bragg scattering and local resonances, which can be engineered to effectively control longitudinal, transverse, and flexural polarized wave propagation. Such characteristics make phononic metamaterials perfect candidates for application in structural elements (e.g., ceramic tiles) to cope with the challenges of noise control in architectures [14, 15, 16, 17, 18]. Numerous studies have explored the potential of phononic crystals for vibration and noise control. For example, Maldovan demonstrated the tunability of acoustic bandgaps in phononic crystals, highlighting their role in controlling low-frequency noise [14]. Similarly, Hussein et al. reviewed the mechanisms of wave attenuation in phononic structures, highlighting their applicability in multi-scale designs for noise control [15]. In the context of architectural applications, research has focused on integrating phononic crystals into building components. Zhao et al. showed how periodic structures embedded in panels could attenuate noise transmission effectively [16], while Ghaffarivardavagh et al. utilized acoustic metamaterials to achieve low-frequency sound attenuation in compact spaces [19]. Another critical aspect of sound control is absorption. Traditional absorptive materials rely on viscous and thermal dissipation mechanisms, but recent studies have shown that metamaterials can achieve efficient sound absorption through localized resonances and impedance matching [20, 21, 22]. For instance, Mei et al. demonstrated that hybrid structures combining phononic crystals and acoustic absorbers could enhance overall performance [20]. Further advancements introduced broadband absorptive designs using subwavelength resonators, paving the way for efficient noise mitigation in compact configurations [22, 23].
Ceramic tile is a widely used construction material which is applied in various parts of buildings, including household roofs, walls, floors, and partitions [24]. In modern architectures, floor tiles are often designed for durability, resistance to abrasion, water impermeability, ease of cleaning, and visual appeal [25]. However, when it comes to acoustic properties, ceramic tiles are not ideal soundproofing materials. Their high density and hard surface make them less effective in reducing sound transmission compared to materials such as wood, laminate, or rubber. Targeting on noise insulation in tiles, Kumar et al. proposed a metatile prototype based on a circular maze-like acoustic metastructure, utilizing a space-coiling technique to enhance acoustic performance in the low-to-mid frequency range [26]. Choi et al. investigated locally resonant metamaterials applied to floors in under-construction residential buildings, conducting an in-depth study on vibro-acoustic control of floor impact sound [27]. Lopes et al. developed a methodology for maximizing broadband impact sound insulation by designing slabs with a non-uniform material thickness distribution, which was numerically optimized [28]. Finally, Sukontasukkul et al. explored the use of viscoelastic polymer sheets as an acoustic control treatment integrated into ceramic tiles to enhance their sound insulation properties [29]. Despite of these significant advancements, achieving effective noise insulation in ceramic tiles remains challenging, particularly in developing an optimal design specifically suited for tiled flooring applications. Previous works have established the role of tile size and material properties in tuning the bandgap frequencies for such applications [30, 16]. Additionally, silicon joints are utilized as coupling elements to optimize wave transmission and reflection, aligning with findings by Fang et al. [30] on the effectiveness of elastomeric joints in acoustic metamaterials.
Building on these principles, this research explores the use of phononic metamaterials and silicon joint design to reduce impact noise in tiled rooms. Inspired by the concept of metatiles [26], a term combining metamaterials and construction tiles, we propose a new design that integrates ceramic plates and silicon. This approach preserves the key properties of ceramic tiles, such as durability, abrasion resistance, water impermeability, ease of cleaning, and aesthetic appeal, while also improving their noise reduction capabilities. First, we analyze and optimize the geometrical parameters and material properties of metatiles, interconnected by silicon joints, to suppress longitudinal and flexural wave propagation. Next, we calculate the dynamic response using the finite element method. Finally, we present experimental validation and evaluate the acoustic performance of ceramic metatiles in impact noise insulation, demonstrating their potential for practical applications in architectural acoustics for tiled floors.
@@
II Theoretical Dispersion and Transmission
II.1 Design of the ceramic metatile
@@
The proposed ceramic metatile is a periodic structure composed of ceramic sub-tiles and silicon joints. Fig. 1(a) shows an isometric view of the main principle behind using ceramic metatiles to mitigate sound propagation from impact noise or footstep noise. As depicted, the metatile floor consists of ceramic metatiles adhered to reinforced concrete using an adhesive. The periodic arrangement of ceramic sub-tiles and silicon joints resembles a monoatomic chain of resonators, with the silicon joints acting as springs and the ceramic sub-tiles serving as masses. This design enables sound wave control through the periodic mass-spring system, using Bragg scattering and local resonance between neighboring sub-tiles. The geometric parameters of the metatile are illustrated in Fig. 1(b), where $a$ represents the periodicity constant of a standard ceramic tile, $c$ is the width of the silicon joints, $b$ is the width of the sub-tiles, and $e$ is the thickness of the ceramic tiles taken as $e=1\,\text{cm}$ in this study. The standard ceramic tiles, provided by IRIS, were characterized through static and dynamic tests under small strains, yielding following estimates for their elastic properties: Young’s modulus $E_{c}=55\pm 5\,\text{GPa}$, Poisson’s ratio $\nu_{c}=0.3$, and density $\rho_{c}=3000\pm 300\,\text{kg/m}^{3}$. For the silicon material, we use a commercially available silicon with the following mechanical properties: Young’s modulus $E_{s}=1\text{MPa}$, Poisson’s ratio $\nu_{s}=0.4$, and density $\rho_{s}=1000\text{kg/m}^{3}$.
II.2 Numerical models
@@
where ${\bf k}=(k,0,0)$ is the reduced Bloch wavevector and $\textbf{u}(x)$ is the displacement vector. The cells are assumed to be periodically repeated along the $x$-direction with periodicity $d=a+5c$.
The phononic dispersion relations are calculated in the first irreducible Brillouin zone $(\Gamma X)$ in the reciprocal space, where $\Gamma=(0,0,0)$ and $X=(\pi/d,0,0)$. For the meshing, we take into account a maximum element size of $\lambda_{\text{min}}/10$ and tetrahedral elements.
Following our examination of the phononic dispersion in ceramic metatiles (infinite), we proceed to investigate their transmission characteristics through simulations involving a single metatile (finite). We compare the results for three different directions for both the standard tile and the metatile, as depicted in Fig. 1(c). The transmission of the elastic waves is calculated using the following equation:
${\rm Transmision}\;{(\rm dB)}=20\log_{10}\left(\frac{\sqrt{|u|^{2}+|v|^{2}+|w|^{2}}_{\rm output}}{\sqrt{|u|^{2}+|v|^{2}+|w|^{2}}_{\rm input}}\right),$ (3)
where $\sqrt{|u|^{2}+|v|^{2}+|w|^{2}}_{\rm input}$ and $\sqrt{|u|^{2}+|v|^{2}+|w|^{2}}_{\rm output}$ are the modulus of the displacement vectors of the input and output elastic waves, respectively.
II.3 Numerical results
The main focus of this study is the investigation of bandgaps, which indicates frequency ranges where the ceramic metatile effectively attenuates footstep and impact noise vibrations. To ensure accurate and reliable results, it is essential to establish a methodology for determining the geometrical dimensions and silicon parameters of the proposed ceramic metatile. We start by analyzing the bandgaps extracted from the phononic dispersion in the first irreducible Brillouin zone, specifically along the $\Gamma X$ direction, for a unit cell consisting of two subtiles of width $b$ and two silicon joints of width $c$. The bandgaps variations are examined as functions of the geometrical parameters of the ceramic metatile: the width of the subtiles $b$ and the width of the silicon joints $c$. Additionally, we investigate the influence of the Young’s modulus of silicon on these bandgaps. The eigenvalue calculations are presented in Fig. 2 for an infinite structure along the $[100]$ direction and $[010]$ direction.
Firstly, we fixed the width of the silicon joints at $c=0.5$ cm and the Young’s modulus of silicon at $1$ MPa. Then, we varied the parameter $b$ from $2$ cm to $10$ cm to analyze its effect on the bandgaps. Fig. 2(a) presents the calculated bandgaps in function of $b$, which are colored in gray, along with their corresponding frequencies. It is observed that for $b=2$ cm, the bandgaps start from $506$ Hz and extend up to approximately $1865$ Hz. As the width of the subtiles increases, the bandgaps shift toward lower frequencies, reaching a minimum frequency of $232$ Hz. This indicates that when the subtiles are larger, bandgaps appear in the lower frequency range. However, as their size continues to increase, the diffraction limit shifts further to lower frequencies, ultimately disrupting the bandgap around $1500$ Hz.

### v2

# Controlling the Propagation of Flexural Elastic Waves With Ceramic Metatiles

@@ Abstract
In this work, we examine the application of phononic metamaterials for elastic impact noise insulation in tiled flooring, through the development of an innovative ceramic metatile that incorporates phononic crystals with optimized joint configurations. First, we optimize the geometrical and material parameters of the proposed metatile, which is composed of small ceramic subtiles connected by silicon joints, in order to reduce longitudinal and flexural wave propagation on tiled floors, which are responsible for noise vibrations in tiled environments. A bandgap is achieved that effectively suppresses the transmission of impact noise through the periodic structural configuration. For flexural waves, the ceramic metatile exhibits a pronounced attenuation of wave transmission in the range of 5005005005001900190019001900 Hz along the $[100]$ direction, and 5005005005001400140014001400 Hz along the $[110]$ direction. For longitudinal waves, a broad bandgap is observed, spanning from $400$ Hz to $1950$ Hz in both the $[100]$ and $[110]$ directions. Additionally, the bandgaps shift toward lower frequencies with increasing width of the subtiles and silicon joints, or with a decrease in the Young’s modulus of the silicon. In both experimental and numerical tests, it is demonstrated that the integration of silicon joints inside the ceramic metatile improves the acoustic insulation performance, as measured by the reduction of impact noise levels across a wide range of low frequencies. The findings highlight the potential of metamaterials in architectural acoustics, offering innovative solutions for elastic wave control in tiled environments.
Brahim Lemkalli1,∗, Ozgur T. Tugut1, Qingxiang Ji1, Richard Craster2,3, Sébastien Guenneau3,4, Muamer Kadic1, Claudio Bizzaglia5, Bogdan Ungureanu3,4,6
{affiliations}
1Université Marie et Louis Pasteur, Institut FEMTO-ST, 25000 Besançon, France
2Mathematics Department, Imperial College London, SW7 2AZ, London, UK
3UMI 2004 Abraham de Moivre-CNRS, Imperial College London, SW7 2AZ, London, UK
4The Blackett Laboratory, Physics Department, Imperial College London, SW7 2AZ, London, UK
5IRIS, Via Guido Reni 2E, 42014 Castellarano, Italy
6METAMAT Limited, 937 Chelsea C, Sloane Avenue, SW3 3EU, London, UK
lemkallibrahim@gmail.com
1 Introduction
@@
One of the most significant factors in building design is to provide a comfortable environment for its residents. Several elements influence a home’s degree of comfort, including temperature, humidity, noise, air quality and hygiene. Neglecting such living conditions will have a negative impact on household comfort, e.g., noise generated by footsteps and noise pollution from surrounding buildings, temperature rise due to heat absorbed by buildings during the daytime, inadequate sanitation, and poor indoor air quality [1, 2, 3]. Among these living conditions, noise control is a critical aspect of architectural design and urban planning, with far-reaching implications for physical and mental health [4, 5]. Reducing noise in enclosed spaces is a long-standing objective in architectural acoustics, especially in contexts like tiled rooms, where vibration reflections and transmissions are enhanced by hard, reflective surfaces [6, 7]. In general, there are numerous ways of controlling noise annoyance both within and outside buildings. Typical noise mitigation methods, such as the use of porous absorbers or heavy insulation, are often encountered with aesthetic, functional, or structural constraints. Recent advances in metamaterials [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], including phononic crystals and acoustic metasurfaces, have opened new avenues for elastic waves control in buildings, enabling the manipulation of wave propagation using compact and tunable designs [6, 22]. Thus, the solution to improve the comfort of noise insulation in building floors should be based on phononic metamaterials in one part and institutive materials in the other part [23, 24, 25].
Phononic metamaterials, composed of periodic arrangements of scatterers and resonators with contrasting acoustic properties, are particularly promising due to their ability to create low-frequency band gaps that inhibit the propagation of specific wave modes [26, 27, 28, 15, 29, 30, 31]. These band gaps arise from mechanisms such as Bragg scattering and local resonances, which can be engineered to effectively control the propagation of longitudinal, transverse, torsional and flexural wave modes [10, 32, 33, 34]. Such capabilities make phononic metamaterials ideal candidates for integration into structural elements (e.g., ceramic tiles) to address noise control challenges in architectural applications [35, 36, 37, 38, 39, 40, 41, 42].
Numerous studies have explored the potential of acoustic metamaterials for vibration and impact noise control [43, 44, 45, 46, 47, 48]. For example, Maldovan demonstrated the tunability of acoustic band gaps in phononic structures, emphasizing their role in mitigating low-frequency noise [35]. Similarly, Hussein et al. reviewed wave attenuation mechanisms in phononic structures, underlining their relevance for multi-scale noise control designs [36]. Ding et al. investigated the effect of unit cell arrangement in a phononic array, introducing the concepts of isotacticity and syndiotacticity to achieve low-frequency and broadband band gaps [49]. Zhou et al. investigated an innovative method for controlling longitudinal wave band gaps, based on the application of an inhomogeneous and periodic thermal field [50]. Wang et al. developed elastic dissipative metasurfaces for the broadband mitigation of Rayleigh waves [51]. In general, the works mentioned achieve band gaps through complex or tunable structures. However, there remains a strong demand for simpler and more adaptable phononic metamaterial that can provide broadband band gaps at low frequencies [52]. Gao et al. achieved two key objectives using phononic metamaterials with time-delayed feedback control. One maximizing the width of the bandgap at a given attenuation rate, and two achieving the highest possible attenuation rate within the targeted frequency range. In architectural contexts, research has increasingly focused on embedding phononic metamaterials into building components. Zhao et al. showed that periodic structures integrated into metamaterial panels can effectively absorb transmitted noise in the 400–800 Hz range [38], while Ghaffarivardavagh et al. employed acoustic metamaterials to achieve low-frequency attenuation in compact spaces, specifically within the 450–500 Hz range [53]. More recently, Lu et al. introduced an ultrathin nonlocal acoustic metaliner, designed using a mode-matching method, which enables global resonance modulation and effectively suppresses anti-resonance between resonance modes, thereby mitigating broadband fan noise. The metaliner operates efficiently in the frequency range of 10005000100050001000–50001000 – 5000 Hz [54]. Xu et al. investigated metamaterial-based absorbers capable of simultaneously absorbing sound and vibration within the peaks’ frequency at $850$ Hz and $1210$ Hz [55]. Zhu et al. studied a multilayer structure designed for low-frequency energy dissipation. They demonstrated that the absorption bandwidth can be broadened by adding more units, for instance, a two-unit multilayer structure achieved absorption in the range of 370 to 520 Hz[56].
@@
Another key aspect of impact noise mitigation is wave absorption. Traditional absorptive materials rely on viscous and thermal dissipation, but recent studies have shown that phononic metamaterials can achieve efficient elastic wave absorption through localized resonances [57, 58, 59]. For instance, Mei et al. demonstrated that hybrid structures combining phononic metamaterials with acoustic absorbers could significantly enhance performance across frequencies from 100 to 1000 Hz [57]. More recent developments have introduced broadband absorptive designs using sub-wavelength resonators, paving the way for compact and effective noise mitigation solutions [59, 60, 61].
Ceramic tile is a widely used construction material applied in various parts of buildings, including household roofs, walls, floors, and partitions [62, 63, 64]. In modern architecture, floor tiles are often designed for durability, resistance to abrasion, water impermeability, ease of cleaning, and aesthetic appeal [65]. However, in terms of acoustic performance, ceramic tiles are not effective at reducing noise or vibrations [66, 66, 67]. Due to their high density and rigid surface, they are less effective at attenuating elastic vibration transmission compared to materials such as wood [68], laminate[69], or rubber [70].
To address the issue of noise insulation in tiled surfaces, several studies have proposed innovative solutions [71]. Kumar et al. developed a metatile prototype based on a circular maze-like acoustic meta-structure, utilizing a space-coiling technique to enhance sound insulation in the low-to-mid frequency range 82016008201600820–1600820 – 1600 Hz. However, this design primarily targets airborne sound rather than impact noise generated by footsteps on floors [72]. Choi et al. investigated locally resonant metamaterials integrated into flooring systems of under-construction residential buildings, focusing on vibro-acoustic control of floor impact noise [73]. Lopes et al. proposed a methodology to maximize broadband impact noise insulation by designing slabs with non-uniform material thickness, optimized through numerical simulations [74]. Additionally, Sukontasukkul et al. explored the integration of viscoelastic polymer sheets into ceramic tiles to enhance noise insulation, particularly at low frequencies. While the performance at low frequencies was similar to conventional tiles, the tiles with viscoelastic layers exhibited superior performance at high frequencies above 1000 Hz [75].
Despite these advancements, achieving effective insulation against impact noise, transmitted via elastic wave propagation, remains a significant challenge in ceramic-tiled floors [76, 77]. Many existing solutions rely on alternative materials or complex designs that may compromise the durability and aesthetic qualities inherent to ceramic tiles. Moreover, some approaches are not easily compatible with conventional tile applications [78]. Previous studies have demonstrated that tile size and material properties play a critical role in tuning bandgap frequencies for noise vibrations insulation applications [79, 80, 81, 38, 82]. In our study, we aim to investigate elastic wave propagation in ceramic-tiled floors, with a particular focus on controlling flexural waves through the application of phononic metamaterials. Additionally, silicon joints are utilized as coupling elements to optimize wave transmission and reflection, aligning with findings by Fang et al. [80] on the effectiveness of elastomeric joints in acoustic metamaterials.
Building on these principles, this research explores the use of phononic metamaterials and silicon joint design to reduce impact elastic waves in tiled rooms. Inspired by the concept of metatiles [72], a term combining metamaterials and construction tiles, we propose a new design that integrates ceramic plates and silicon. This approach preserves the key properties of ceramic tiles, such as durability, abrasion resistance, water impermeability, ease of cleaning, and aesthetic appeal, while also improving their elastic waves reduction capabilities. First, we analyze and optimize the geometrical parameters and material properties of metatiles, interconnected by silicon joints, to suppress longitudinal and flexural wave propagation. Next, we calculate the dynamic response using the finite element method. Finally, we present experimental validation and evaluate the acoustic performance of ceramic metatiles in impact elastic waves insulation, demonstrating their potential for practical applications in architectural acoustics for tiled floors. The results reveal broadband band gaps that cover the frequency range typically generated by footsteps and indoor noise. To characterize our design, we have structured the paper as follows: We begin by describing the design of the ceramic metatile, followed by the optimization of its geometrical parameters. We then present numerical results for two types of elastic wave polarizations, flexural and longitudinal, as detailed in Section 2. Next, we describe the laboratory setup used to experimentally measure the transmission of flexural waves, along with the data processing performed using the Fast Fourier Transform (FFT) in Section 2. Finally, we examine a real-world scenario involving the generation of elastic impact noise. The corresponding FFT analysis is presented in Section 3.2, demonstrating the effectiveness of our ceramic metatile-based design.
2 Theoretical Dispersion and Transmission
This section presents the design methodology of the ceramic metatiles and the optimization of the geometrical parameters of their constituent units. Additionally, it describes the numerical methods used to calculate transmission for both longitudinal and flexural waves.
2.1 Design of the ceramic metatile
@@
The proposed ceramic metatile is a periodic structure composed of ceramic sub-tiles and silicon joints. Figure 1(a) shows an isometric view of the main principle behind using ceramic metatiles to mitigate elastic wave propagation from impact noise or footstep noise. As depicted, the metatile floor consists of ceramic metatiles adhered to reinforced concrete using an adhesive. The periodic arrangement of ceramic sub-tiles and silicon joints resembles a monoatomic chain of resonators, with the silicon joints acting as springs and the ceramic sub-tiles serving as masses. This design enables elastic wave control through the periodic mass-spring system, using Bragg scattering and local resonance between neighboring sub-tiles. The geometric parameters of the metatile are illustrated in Figure 1(b), where $a$ represents the periodicity constant of a standard ceramic tile, $c$ is the width of the silicon joints, $b$ is the width of the subtiles, and $e$ is the thickness of the ceramic tiles taken as $e=1\,\text{cm}$ in this study. The standard ceramic tiles, provided by IRIS, were characterized through static and dynamic tests under small strains, yielding following estimates for their elastic properties: Young’s modulus $E_{c}=55\pm 5\,\text{GPa}$, Poisson’s ratio $\nu_{c}=0.3$, and density $\rho_{c}=3000\pm 300\,\text{kg/m}^{3}$. For the silicon material, we use commercially available silicon with the following mechanical properties: Young’s modulus $E_{s}=1\text{MPa}$, Poisson’s ratio $\nu_{s}=0.4$, and density $\rho_{s}=1000\text{kg/m}^{3}$.
2.2 Numerical models
@@
where ${\bf k}=(k,0,0)$ is the reduced Bloch wavevector and $\textbf{u}(x)$ the displacement vector at point $x$. Cells are assumed to be periodically repeated along the $x$-direction with periodicity $d=a+5c$.
The phononic dispersion relations are calculated in the first irreducible Brillouin zone $(\Gamma X)$ in the reciprocal space, where $\Gamma=(0,0,0)$ and $X=(\pi/d,0,0)$. For the meshing, we take into account a maximum element size of $\lambda_{\text{min}}/10$ and tetrahedral elements. To delineate the flexural waves, we computed the dispersion curves, taking into account the polarization weighting using the following expression:
@@
where $V$ is the volume of the elementary cell $|u|=\sqrt{uu^{*}}$ is the modulus of displacement vector, with $u^{*}$ the complex conjugate of $u$.
Following our examination of the phononic dispersion in ceramic metatiles (infinite), we proceed to investigate their transmission characteristics through simulations involving a single metatile (finite). We compare the results for three different directions for both the standard tile and the metatile, as depicted in Figure 1(c). The transmission of the elastic waves is calculated using the following equation:
${\rm Transmision}\;{(\rm dB)}=20\log_{10}\left(\frac{|u_{i}|_{\rm output}}{|u_{i}|_{\rm input}}\right),$ (4)
where $|u_{i}|_{\text{input}}$ and $|u_{i}|_{\text{output}}$ denote the magnitude of the displacement vector component (i.e., $\sqrt{uu^{*}}$, $\sqrt{vv^{*}}$, or $\sqrt{ww^{*}}$) at the input and output of the ceramic metatile, respectively. The component $u_{i}\in\{u,v,w\}$ is selected based on the direction of the input wave and depends on the nature of the wave polarization: longitudinal or flexural.
2.3 Numerical results
The main focus of this study is the investigation of bandgaps, which indicates frequency ranges where the ceramic metatile effectively attenuates footstep and impact noise vibrations. To get accurate and trustworthy results, it is important to create a clear method for improving the size and material features of the proposed ceramic subtile and its silicon joints. To achieve this, we conduct an optimization study aimed at obtaining broadband band gaps and determining the optimal geometrical and material parameters. The study is performed using COMSOL Multiphysics, with manual optimization carried out through a parametric sweep approach; one parameter is fixed while the other is varied. We begin by varying the width of the subtiles while keeping the other parameters constant to identify the optimal width. Once the optimal width is determined, it is fixed, and we proceed to vary the thickness of the silicon joints. Finally, we study the influence of the material properties of the joints by performing a parametric sweep on their mechanical characteristics by analyzing the bandgaps extracted from the phononic dispersion in the first irreducible Brillouin zone, specifically along the $\Gamma X$ direction, for a unit cell consisting of two subtiles of width $b$ and two silicon joints of width $c$.
The bandgaps variations are examined as functions of the geometrical parameters of the ceramic metatile: the width of the subtiles $b$ and the width of the silicon joints $c$. Additionally, we investigate the influence of the Young’s mosdulus of silicon on these bandgaps. The eigenvalue calculations are presented in Figure 2 for an infinite structure along the $[100]$ and $[010]$ crystallographic directions.
Firstly, we fixed the width of the silicon joints at $c=0.5$ cm and the Young’s modulus of silicon at $1$ MPa. Then, we varied the parameter $b$ from $2$ cm to $10$ cm to analyze its effect on the bandgaps. Figure 2(a) presents the calculated bandgaps in function of $b$, which are colored in gray, along with their corresponding frequencies. It is observed that for $b=2$ cm, the bandgaps start from $506$ Hz and extend up to approximately $1865$ Hz. As the width of the sub-tiles increases, the bandgaps shift toward lower frequencies, reaching a minimum of $232$ Hz. This behavior is consistent with the classical mass-spring system, where larger masses result in lower resonance frequencies due to the inverse relationship between frequency and mass. In such systems, both the mass and the stiffness of the spring play a critical role in determining the resonance frequency. Additionally, the increase in the sub-tile width $b$ influences Bragg scattering, as it alters the periodicity of the structure. In this case, the wavelength associated with Bragg scattering is approximately $\lambda=2\times b$; thus, as $b$ increases, the wavelength increases, leading to a corresponding decrease in frequency. This explains the observed downward shift in the bandgap frequencies with increasing $b$. This indicates that when the subtiles are larger, bandgaps appear in the lower frequency range. However, as their size continues to increase, the diffraction limit shifts further to lower frequencies, ultimately disrupting the bandgap around $1500$ Hz.
[…]